A stone is thrown the top of building upward at an angle of 30 to the horizontal with an initial speed of 20 m/s , if the height of the building is 45 m. How long does the stone to reach the ground? What is the speed of the stone just before it strikes ground?
1
Expert's answer
2020-05-08T16:21:34-0400
"y=h+v\sin{30}t-gt^2=0\\ 45+20\sin{30}t-9.8t^2=0"

"t=2.7\ s"